# Intertwined order of generalized global symmetries

### v1

@@ Abstract
We investigate the interplay of generalized global symmetries in 2+1 dimensions by introducing a lattice model that couples a $\mathbb{Z}_{N}$ clock model to a $\mathbb{Z}_{N}$ gauge theory via a topological interaction. This coupling binds the charges of one symmetry to the disorder operators of the other, and when these composite objects condense, they give rise to emergent generalized symmetries with mixed ’t Hooft anomalies. These anomalies result in phases with ordinary symmetry breaking, topological order, and symmetry-protected topological (SPT) order, where the different types of order are not independent but intimately related. We further explore the gapped boundary states of these exotic phases and develop theories for phase transitions between them. Additionally, we extend our lattice model to incorporate a non-invertible global symmetry, which can be spontaneously broken, leading to domain walls with non-trivial fusion rules.
@@ 1 Introduction
Symmetry plays an fundamental role in physics and particularly in condensed matter physics. In the conventional Landau theory of phase transitions, phases of matter are classified by the global symmetries they spontaneously break. On the other hand, symmetry-protected topological phases (SPTs) are characterized by their nontrivial response to probes of global symmetries or by the ’t Hooft anomalies of their boundary states [1, 2, 3, 4, 5, 6].
@@
Recent developments have expanded the concept of global symmetry in several ways, thus widening our perspective on how to characterize phases of matter [7, 8, 9, 10]. While an ordinary global symmetry acts on local operators, a $q$-form global symmetry acts on operators supported on $q$-dimensional objects. Using this terminology, an ordinary global symmetry is a zero-form symmetry. We will be especially interested in one-form global symmetries, which act on operators defined on noncontractible loops, since these symmetries can be used to characterize topologically ordered phases [11, 12]. Indeed, the braiding of two Abelian anyons in (2+1)d results in a phase factor, which can be interpreted as resulting from the action of a one-form symmetry by the worldline of one anyon on another. A phase with Abelian topological order may then be thought of as “spontaneously” breaking a discrete one-form global symmetry in the sense that the symmetry is respected by the low energy effective action (or Hamiltonian) but not by a given ground state [^1] For ordinary zero-form global symmetries, the precise definition of spontaneous symmetry breaking is that upon coupling to a local external symmetry-breaking field $h(x)$, the order parameter remains nonzero if we take the thermodynamic limit and then $h\to 0$. A analogue of this criterion does not exist for higher-form symmetries because the observables are non-local. However, a spontaneously broken one-form global symmetry can be defined as the condition that the correlators of Wilson operators on noncontractible loops obey clustering. A well-known example is the Polyakov loop used to diagnose the thermal phase transition from a confined to a deconfined phase of a pure gauge theory [13].. We emphasize that this one-form symmetry is an emergent infrared (IR) symmetry in the usual condensed matter setting, but once it emerges, it is extremely robust [14].
@@
A distinct generalization of symmetry that has garnered much attention recently is non-invertible symmetry [15, 16]. Conventional symmetry operators are unitary (or antiunitary) and thus have inverses. Recently, it has been argued that topological operators that do not necessarily have inverses should be viewed as symmetry operators [17, 18]. While ordinary global symmetries are characterized mathematically by groups, symmetry operators for non-invertible symmetries can have multiple fusion channels. A flurry of recent work has introduced many examples of field theories [19, 20, 21, 22] and lattice models [23, 24, 25] with non-invertible symmetries, demonstrating that these symmetries can lead to novel SPTs [26, 27] and intricate phase diagrams with exotic symmetry-breaking patterns [28, 29, 30].
@@
Experience with ordinary global symmetries suggests that the interplay of different symmetries can give rise to rich phase diagrams with a variety of interesting phases. In particular, the concept of intertwined order has been invoked to understand general principles at play in the complex phase diagrams of correlated electronic systems [31]. The idea is that the different orders in a complex phase diagram can arise from an underlying microscopic primary state. A paradigmatic example of a possible parent state is a pair density wave superconductor [32], which spontaneously breaks $U(1)$ particle number conservation and translation symmetry, which hosts several ordered states.
@@
We therefore should expect that a rich set of physics to arise from models with an interplay of generalized global symmetries. In this work, we develop such models with intertwined generalized symmetries. One way in which generalized symmetries can be intertwined is that they can form a higher-group symmetry [33, 34], but we take a different approach here. With some inspiration from a previous work of the authors [35], our starting point is a 3d Euclidean lattice model consisting of a $\mathbb{Z}_{N}$ clock model coupled to a $\mathbb{Z}_{N}$ gauge theory (defined on dual lattices) by a topological term rather than by minimal coupling [^2] In this model the $\mathbb{Z}_{N}$ global symmetry of the clock model is not gauged and the $\mathbb{Z}_{N}$ local symmetry of the gauge theory is not Higgsed.. This topological coupling is analogous to the theta term in (3+1)d gauge theories, which gives magnetic monopoles an electric charge—a phenomenon known as the Witten effect [36]. Similarly, our (2+1)d topological interaction binds vortices of the spin model to an electric charge of the gauge theory, and the magnetic monopoles of the gauge theory become bound to the spins. Different combinations of these composite operators can condense, leading to a rich phase diagram with a variety of interesting phases. Using duality arguments, we can deduce much of this phase diagram and the physics of the phases in the model.
@@
where $A_{\mu}$ is a $\mathbb{Z}_{N}$ one-form background gauge field, $B_{\mu\nu}$ is a $\mathbb{Z}_{N}$ two-form background gauge field, and $p\in\mathbb{Z}$ is a parameter characterizing the SPT phase. A lattice model for an SPT of this type was first described for $N=2$ and $p=1$ on the triangular lattice in Ref. [37]. For a given $N$, an SPT with each possible value of $p$ is realized in a phase of the lattice model. Our lattice model thus serves as a microscopic model for every SPT phase with a response of the form in Eq. (1.1), thus providing a versatile platform to study these phases and their phase transitions.
@@
Our lattice model also admits more intricate phases characterized by an effective field theory that is a gauged version of Eq. (1.1), meaning that $A_{\mu}$ and $B_{\mu\nu}$ are both promoted to fluctuating gauge fields. By the correspondence between SPTs and topologically ordered states [38], these phases are symmetry enriched topological states (SETs) that have topological order, but there is some additional symmetry breaking of the ordinary global symmetry. These phases are generated by simultaneously condensing bound states of local operators with spin charge $N$ and magnetic charge $p$ and loop operators with electric charge $N$ and vorticity $p$, which is reminiscent of the physics of oblique confining phases in (3+1)d gauge theories [39, 40, 41], so we refer to this class of (2+1)d phases as oblique phases. We demonstrate that for $L=\gcd(N,p)>1$, the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry has mixed ’t Hooft anomalies with emergent generalized symmetries, resulting in a breaking of this symmetry to a nontrivial subgroup, $H=\mathbb{Z}_{N/L}^{(0)}\times\mathbb{Z}_{N/L}^{(1)}$, giving rise to both ordinary symmetry breaking and topological order. Furthermore, the unbroken subgroup $H$ has a nontrivial response to background fields, signaling SPT order for this symmetry.
@@
To emphasize the new features of our lattice model, let us contrast with another physical system that has both ordinary symmetry breaking and topological order. For example, fractional quantum Hall systems can certainly host broken symmetry states with nematic order that coexists with topological order [42, 43]. However, the symmetry structure of the systems we are studying here are more intricate. The physics described above for an oblique phase is representative of the interplay of the zero-form and one-form symmetries in our lattice model more generally. The patterns of symmetry breaking for the two symmetries are not independent and do not merely coexist. For any phase of our lattice model, if the $\mathbb{Z}_{N}$ zero-form symmetry is spontaneously broken to a nontrivial subgroup, the $\mathbb{Z}_{N}$ one-form symmetry is also broken to the same subgroup. The remaining unbroken zero-form and one-form symmetries additionally have mixed SPT order. These features demonstrate that the symmetries are not independent but indeed have a special interplay throughout the phase diagram because of the way in which they are coupled in the lattice model.
@@
For systems with an open boundary, for every SPT of the form in Eq. (1.1), preserving the zero-form and one-form symmetries inevitably leads to boundary modes. The one-form symmetry cannot be spontaneously broken along the (1+1)d boundary, so any gapped boundary state must necessarily break the zero-form symmetry spontaneously. We develop the precise criteria the boundary theory must satisfy to cancel the ’t Hooft anomaly of the bulk and provide examples. We also develop two different gapped boundary states for oblique phases. One of these boundary conditions, which we refer to as an electric boundary condition, preserves the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry of the bulk. This boundary condition is consistent with the SPT order of the unbroken $H=\mathbb{Z}_{N/L}^{(0)}\times\mathbb{Z}_{N/L}^{(1)}$ subgroup and results in the spontaneous breaking of $H$ at the boundary. The other gapped boundary condition, which we call the magnetic boundary state, preserves a magnetic $\mathbb{Z}_{p}^{(0)}\times\mathbb{Z}_{p}^{(1)}$ symmetry, resulting in the spontaneous breaking of this symmetry at the boundary.
@@
We also find possible gapless boundary states for $\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ SPTs. By analyzing possible perturbations that gap the boundary, we show that the gapless state can be interpreted as a quantum critical point (or line) that has additional emergent symmetries. In the simplest case, we consider the gapless boundary state of the SPT, Eq. (1.1), and add only perturbations that preserve the $\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry coupled to the bulk. For $L=\gcd(N,p)\geq 4$, our gapless state is a quantum critical point (critical line for $L>4$) where the global symmetry is enhanced to $U(1)^{(0)}\times U(1)^{(0)}\times\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$. This critical point or line then separates two gapped phases, each of which has a different $\mathbb{Z}_{N}^{(0)}$ symmetry that is spontaneously broken.
@@
If we now gauge the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry of the lattice model, the domain walls will now separate distinct oblique states characterized by different topological orders, spontaneous symmetry breaking, and mixed SPT orders. As with the boundary states of the oblique phases, we must then necessarily decorate the domain walls with dynamical degrees of freedom, which is somewhat reminiscent of the decorated domain wall construction for SPTs [44], but here, the domain walls are decorated with a lower-dimensional state with spontaneous symmetry breaking rather than SPT order. The physical consequence of introducing these degrees of freedom is that the domain walls no longer obey group-like fusion rules. This generalization of our lattice model thus has a phase in which a non-invertible symmetry is spontaneously broken. Indeed, the mixed anomaly implies that gauging $G$ anomalously breaks the $(\mathbb{Z}_{N}^{(0)})_{\mathrm{axion}}$ symmetry, but we demonstrate that it can alternatively be viewed as a non-invertible symmetry. This notion of generating non-invertible symmetries from mixed anomalies as appeared in other contexts previously [19, 21].
This paper is organized as follows. In Section 2, we introduce our lattice model and discuss its global symmetries and phase diagram. Section 3 describes the low energy physics and effective field theory of the oblique phases. In Section 4, we discuss the intertwined SPT response of a generic oblique phase. Section 5 develops the phase transitions of the bulk phases of the lattice model in more detail. We then derive possible gapped boundary states of both the SPT state and the oblique phases in Section 6. Section 7 introduces a Hamiltonian lattice model that realizes oblique phases as its ground states. In Section 8, we develop gapless boundary states of the SPT states and present examples of boundary phase transitions into other gapped boundary states. Finally, in Section 9, we study the two $\mathbb{Z}_{N}$ axion generalizations of our lattice model, one of which has mixed anomalies of generalized symmetries in the UV, and another which has a non-invertible symmetry. In Appendix A, we review $\mathbb{Z}_{N}$ gauge fields in the continuum. In Appendix B, we find dualities of our lattice model. In Appendix C, we derive an effective Coulomb gas action for our lattice model and relate it to duality. Appendix D reviews the lattice version of (2+1)d BF theory and its counterpart in the ordered phase of the clock model. Appendix E contains some of the more technical calculations associated with the $\mathbb{Z}_{N}$ axion models.

### v2

@@ Abstract
We investigate the interplay of generalized global symmetries in 2+1 dimensions in a lattice model that couples a $\mathbb{Z}_{N}$ clock model to a $\mathbb{Z}_{N}$ gauge theory via a topological interaction. This coupling binds the charges of one symmetry to the disorder operators of the other, and when these composite objects condense, they give rise to emergent generalized symmetries with mixed ’t Hooft anomalies. These anomalies result in phases with ordinary symmetry breaking, topological order, and symmetry-protected topological (SPT) order, where the different types of order are not independent but intimately related. We further explore the gapped boundary states of these exotic phases and develop theories for phase transitions between them. Additionally, we extend this lattice model to incorporate a non-invertible global symmetry, which can be spontaneously broken, leading to domain walls with non-trivial fusion rules.
@@ 1 Introduction
Symmetry plays a fundamental role in physics and particularly in condensed matter physics. In the conventional Landau theory of phase transitions, phases of matter are classified by the global symmetries they spontaneously break. Furthermore, symmetry-protected topological phases (SPTs) are characterized by their nontrivial response to probes of global symmetries or by the ’t Hooft anomalies of their boundary states [1, 2, 3, 4, 5, 6].
@@
Recent developments have expanded the concept of global symmetry in several ways, thus widening our perspective on how to characterize phases of matter [7, 8, 9, 10]. While an ordinary global symmetry acts on local operators, a $q$-form global symmetry acts on operators supported on $q$-dimensional objects. Using this terminology, an ordinary global symmetry is a zero-form symmetry. We will be especially interested in one-form global symmetries, which act on operators defined on noncontractible loops, since these symmetries can be used to characterize topologically ordered phases [11, 12]. Indeed, the braiding of two Abelian anyons in (2+1)d results in a phase factor, which can be interpreted as resulting from the action of a one-form symmetry by the worldline of one anyon on another. A phase with Abelian topological order may then be thought of as “spontaneously” breaking a discrete one-form global symmetry in the sense that the symmetry is respected by the low energy effective action (or Hamiltonian) but not by a given ground state [^1] For ordinary zero-form global symmetries, the precise definition of spontaneous symmetry breaking is that upon coupling to a local external symmetry-breaking field $h(x)$, the order parameter remains nonzero if we take the thermodynamic limit and then $h\to 0$. A analogue of this criterion does not exist for higher-form symmetries because the observables are non-local. However, a spontaneously broken one-form global symmetry can be defined as the condition that the correlators of Wilson operators on noncontractible loops obey clustering. A well-known example is the Polyakov loop used to diagnose the thermal phase transition from a confined to a deconfined phase of a pure gauge theory [13].. We emphasize that this one-form symmetry is an emergent infrared (IR) symmetry in the usual condensed matter setting, but once it emerges, it is highly robust [14].
@@
A distinct generalization of symmetry that has garnered much attention recently is non-invertible symmetry [15, 16]. Conventional symmetry operators are unitary (or antiunitary) and thus have inverses. Recently, it has been argued that topological operators that do not necessarily have inverses should be viewed as symmetry operators [17, 18]. While ordinary global symmetries are characterized mathematically by groups, symmetry operators for non-invertible symmetries can have multiple fusion channels. A flurry of recent work has introduced many examples of field theories [19, 20, 21, 22] and lattice models [23, 24, 25, 26, 27] with non-invertible symmetries, demonstrating that these symmetries can lead to novel SPTs [28, 29] and intricate phase diagrams with exotic symmetry-breaking patterns [30, 31, 32].
@@
Experience with ordinary global symmetries suggests that the interplay of different symmetries can give rise to rich phase diagrams with a variety of interesting phases. In particular, the concept of intertwined order was introduced to understand general principles at play in the complex phase diagrams of correlated electronic systems [33]. The idea is that the different orders in a complex phase diagram can arise from an underlying microscopic primary state. A paradigmatic example of a possible parent state is a pair density wave superconductor [34], which spontaneously breaks $U(1)$ particle number conservation and translation symmetry, hosting several ordered states.
@@
We therefore should expect a rich set of physics to arise from models with an interplay of generalized global symmetries. In this work, we explore models with intertwined generalized symmetries. One way in which generalized symmetries can be intertwined is that they can form a higher-group symmetry [35, 36], but we take a different approach here. With some inspiration from a previous work of the authors [37], our starting point is a 3d Euclidean lattice model consisting of a $\mathbb{Z}_{N}$ clock model coupled to a $\mathbb{Z}_{N}$ gauge theory (defined on dual lattices) by a topological term rather than by minimal coupling [^2] In this model the $\mathbb{Z}_{N}$ global symmetry of the clock model is not gauged and the $\mathbb{Z}_{N}$ local symmetry of the gauge theory is not Higgsed.. A more general self-dual lattice model, which includes the present one, was first introduced by Shapere and Wilczek [38], who demonstrated that the model is self-dual, though they did not explore its global symmetries or the topological physics of its phases as we do here. The topological coupling in the model is analogous to the theta term in (3+1)d gauge theories, which gives magnetic monopoles an electric charge—a phenomenon known as the Witten effect [39]. Similarly, this (2+1)d interaction binds vortices of the spin model to an electric charge of the gauge theory, and the magnetic monopoles of the gauge theory become bound to the spins. Different combinations of these composite operators can condense, leading to a rich phase diagram with a variety of interesting phases. Using duality arguments, we can deduce much of this phase diagram and the physics of the phases in the model.
@@
where $A_{\mu}$ is a $\mathbb{Z}_{N}$ one-form background gauge field, $B_{\mu\nu}$ is a $\mathbb{Z}_{N}$ two-form background gauge field, and $p\in\mathbb{Z}$ is a parameter characterizing the SPT phase. A lattice model for an SPT of this type was first described for $N=2$ and $p=1$ on the triangular lattice in Ref. [40]. For a given $N$, an SPT with each possible value of $p$ is realized in a phase of the lattice model we study. This lattice model thus serves as a microscopic model for every SPT phase with a response of the form in Eq. (1.1), thus providing a versatile platform to study these phases and their phase transitions.
@@
This lattice model also admits more intricate phases characterized by an effective field theory that is a gauged version of Eq. (1.1), meaning that $A_{\mu}$ and $B_{\mu\nu}$ are both promoted to fluctuating gauge fields. By the correspondence between SPTs and topologically ordered states [41], these phases are symmetry enriched topological states (SETs) that have topological order, but there is some additional symmetry breaking of the ordinary global symmetry. These phases are generated by simultaneously condensing bound states of local operators with spin charge $N$ and magnetic charge $p$ and loop operators with electric charge $N$ and vorticity $p$, which is reminiscent of the physics of oblique confining phases in (3+1)d gauge theories [42, 43, 44], so we refer to this class of (2+1)d phases as oblique phases. We demonstrate that for $L=\gcd(N,p)>1$, the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry has mixed ’t Hooft anomalies with emergent generalized symmetries, resulting in the breaking of this symmetry to a nontrivial subgroup, $H=\mathbb{Z}_{N/L}^{(0)}\times\mathbb{Z}_{N/L}^{(1)}$, giving rise to both ordinary symmetry breaking and topological order. Furthermore, the unbroken subgroup $H$ has a nontrivial response to background fields, signaling SPT order for this symmetry.
@@
To emphasize the new features within this lattice model, let us contrast with another physical system that has both ordinary symmetry breaking and topological order. For example, fractional quantum Hall systems can certainly host broken symmetry states with nematic order that coexists with topological order [45, 46]. However, the symmetry structure of the systems we are studying here are more intricate. The physics described above for an oblique phase is representative of the interplay of the zero-form and one-form symmetries in our lattice model more generally. The patterns of symmetry breaking for the two symmetries are not independent and do not merely coexist. For any phase of our lattice model, if the $\mathbb{Z}_{N}$ zero-form symmetry is spontaneously broken to a nontrivial subgroup, the $\mathbb{Z}_{N}$ one-form symmetry is also broken to the same subgroup. The remaining unbroken zero-form and one-form symmetries additionally have mixed SPT order. These features demonstrate that the symmetries are not independent but indeed have a special interplay throughout the phase diagram because of the way in which they are coupled in the lattice model.
@@
For systems with an open boundary, for every SPT of the form in Eq. (1.1), preserving the zero-form and one-form symmetries inevitably leads to boundary modes. The one-form symmetry cannot be spontaneously broken along the (1+1)d boundary [^3] Since we are dealing with a discrete one-form symmetry, this fact is equivalent to the statement that there is no topological order in (1+1)d. [8, 47], so any gapped boundary state must necessarily break the zero-form symmetry spontaneously. We develop the precise criteria that the boundary theory must satisfy to cancel the ’t Hooft anomaly of the bulk and provide examples. We also develop two different gapped boundary states for oblique phases. One of these boundary conditions, which we refer to as an electric boundary condition, preserves the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry of the bulk. This boundary condition is consistent with the SPT order of the unbroken $H=\mathbb{Z}_{N/L}^{(0)}\times\mathbb{Z}_{N/L}^{(1)}$ subgroup and results in the spontaneous breaking of $H$ at the boundary. The other gapped boundary condition, which we call the magnetic boundary state, preserves a magnetic $\mathbb{Z}_{p}^{(0)}\times\mathbb{Z}_{p}^{(1)}$ symmetry, resulting in the spontaneous breaking of this symmetry at the boundary.
@@
We also find possible gapless boundary states for $\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ SPTs. By analyzing possible perturbations that gap the boundary, we show that the gapless state can be interpreted as a quantum critical point (or critical line) that has additional emergent symmetries. In the simplest case, we consider the gapless boundary state of the SPT, Eq. (1.1), and add only perturbations that preserve the $\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry coupled to the bulk. For $L=\gcd(N,p)\geq 4$, our gapless state is a quantum critical point (critical line for $L>4$) where the global symmetry is enhanced to $U(1)^{(0)}\times U(1)^{(0)}\times\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$. This critical point or line then separates two gapped phases, each of which has a different $\mathbb{Z}_{N}^{(0)}$ symmetry that is spontaneously broken.
@@
If we now gauge the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry of the lattice model, the domain walls will separate distinct oblique states characterized by different topological orders, spontaneous symmetry breaking, and mixed SPT orders. As with the boundary states of the oblique phases, we must then necessarily decorate the domain walls with dynamical degrees of freedom, which is somewhat reminiscent of the decorated domain wall construction for SPTs [48], but here, the domain walls are decorated with a lower-dimensional state with spontaneous symmetry breaking rather than SPT order. The physical consequence of introducing these degrees of freedom is that the domain walls no longer obey group-like fusion rules. This generalization of our lattice model thus has a phase in which a non-invertible symmetry is spontaneously broken. Indeed, the mixed anomaly implies that gauging $G$ anomalously breaks the $(\mathbb{Z}_{N}^{(0)})_{\mathrm{axion}}$ symmetry, but we demonstrate that it can alternatively be viewed as a non-invertible symmetry. This notion of generating non-invertible symmetries from mixed anomalies has appeared in other contexts previously [19, 21].
This paper is organized as follows. In Section 2, we introduce the 3d lattice model and discuss its global symmetries and phase diagram. Section 3 describes the low energy physics and effective field theory of the oblique phases. In Section 4, we discuss the intertwined SPT response of an oblique phase. Section 5 develops the phase transitions of the bulk phases of the lattice model in more detail. We then derive possible gapped boundary states of both the SPT state and the oblique phases in Section 6. Section 7 introduces a Hamiltonian lattice model that realizes oblique phases as its ground states. In Section 8, we develop gapless boundary states of the SPT states and present examples of boundary phase transitions into other gapped boundary states. Finally, in Section 9, we study the two $\mathbb{Z}_{N}$ axion generalizations of our lattice model, one of which has mixed anomalies of generalized symmetries in the UV, and another which has a non-invertible symmetry. In Appendix A, we review $\mathbb{Z}_{N}$ gauge fields in the continuum. In Appendix B, we find dualities of our lattice model. In Appendix C, we derive an effective Coulomb gas action for our lattice model and relate it to duality. Appendix D reviews the lattice version of (2+1)d BF theory and its counterpart in the ordered phase of the clock model. Appendix E contains some of the more technical calculations associated with the $\mathbb{Z}_{N}$ axion models.